OpenStudy (anonymous):

how do i solve (2x+5)/(x^2-3x-10) + (x+1)/(x+2)

OpenStudy (muzzack):

x • (x - 2) ————————————————— (x + 2) • (x - 5)

OpenStudy (muzzack):

want me to do it step by stpe

OpenStudy (solomonzelman):

Hint, \(\normalsize\color{blue}{ \rm x^2-3x-10=(x+2)(x-5) }\) now, find the common denominator.

OpenStudy (anonymous):

\[\frac{ 2x+5 }{ x ^{2}-3x-10 } + \frac{ x+1 }{ x+2 }\]

OpenStudy (anonymous):

yes please @Muzzack

OpenStudy (muzzack):

Step 1: You will have to simplify \[\frac{2x + 5 }{ x^2 - 3x - 10 } \]

OpenStudy (muzzack):

then you get (2x + 5) (x + 1) ————————————————— + ——————— (x + 2) • (x - 5) x + 2

OpenStudy (anonymous):

ok then what?

OpenStudy (muzzack):

Step 2 : 2x+5 x+1 Simplify ——————————— + ————— (x+2)•(x-5) (x+2)

OpenStudy (muzzack):

Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: (2x+5) + (x+1) • (x-5) x^2 - 2x —————————————————————— = ————————————————— (x+2) • (x-5) (x + 2) • (x - 5) Pull out like factors : x^2 - 2x = x • (x - 2) Final result : x • (x - 2) ————————————————— (x + 2) • (x - 5)

OpenStudy (anonymous):

oh my gosh thanks I get it now!

OpenStudy (muzzack):

ur welcome